A 0.2025g sample consisting of only BaCl2 and KCl required 20.25ml of 0.1200M AgNO3 solution for the quantitative precipitation of chloride. Calculate the %Ba and %K in the sample


1
Expert's answer
2021-02-09T04:37:36-0500

Solution:

let x = mass of KCl, y = mass of BaCl2

Thus: x + y = 0.2025 g; → x = 0.2025 - y


The molar mass of BaCl2 is 208.23 g mol-1

The molar mass of KCl is 74.55 g mol-1


Moles of AgNO3 = Molarity of AgNO3 × Solution volume

Moles of AgNO3 = 0.1200 M × 0.02025 L = 0.00243 mol

Moles of AgNO3 = 0.00243 mol


Balanced chemical equations:

KCl + AgNO3 → AgCl + KNO3

BaCl2 + 2AgNO3 → 2AgCl + Ba(NO3)2

According to these equations above:

Moles AgNO3 from KCl = (x g KCl × 1 mol AgNO3) / (74.55 g mol-1 KCl) = 0.01341x mol AgNO3

Moles AgNO3 from BaCl2 = (y g BaCl2 × 2 mol AgNO3) / (208.23 g mol-1 BaCl2l) = 0.00960y mol AgNO3


0.01341x + 0.00960y = 0.00243

0.01341 × (0.2025 - y) + 0.00960y = 0.00243

0.002716 - 0.01341y + 0.00960y = 0.00243

-0.0038y = -0.00027

y = 0.07506 g - mass of BaCl2

x = 0.2025 - y = 0.2025 - 0.07506 = 0.12744

x = 0.12744 g - mass of KCl


Moles of BaCl2 = Mass of BaCl2 / Molar mass of BaCl2

Moles of of BaCl2 = 0.07506 g / 208.23 g mol-1 = 0.0003605 mol BaCl2

BaCl2 → Ba2+ + 2Cl-

Hence,

Moles of Ba2+ = Moles of BaCl2 = 0.0003605 mol

Mass of Ba2+ = Moles of Ba2+ × Molar mass of Ba2+

The molar mass of Ba is 137.33 g mol-1

Hence,

Mass of Ba2+ = 0.0003605 mol × 137.33 g mol-1 = 0.04951 g

%Ba = (0.04951 g / 0.2025 g) × 100% = 24.45%

%Ba = 24.45%


Moles of KCl = Mass of KCl / Molar mass of KCl

Moles of of KCl = 0.12744 g / 74.55 g mol-1 = 0.001709 mol KCl

KCl → K+ + Cl-

Hence,

Moles of K+ = Moles of KCl = 0.001709 mol

Mass of K+ = Moles of K+ × Molar mass of K+

The molar mass of K is 39.098 g mol-1

Hence,

Mass of K+ = 0.001709 mo × 39.098 g mol-1 = 0.06682 g

%K = (0.06682 g / 0.2025 g) × 100% = 32.99% = 33.00%

%K = 33.00%


Answer:

%Ba = 24.45%;

%K = 33.00%.